An infinitely long conductor has a circular loop tangent to it. A current of 5 A flows through the conductor. Find the radius of the loop if the intensity of the magnetic field at its centre is 41 A/m.



1
Expert's answer
2022-01-30T15:27:26-0500

We know that

"B=\frac{\mu_0 ni}{2r}"

Now "B=\mu_0H"

Both equation equal

"\mu_0H=\frac{\mu_0 ni}{2r}"

"H=\frac{ni}{2r}"

Put n=1

i=5A

H=41A/m

Put value

"41=\frac{1\times5}{2\times r}"

"r=\frac{5}{82}=0.060m"